OpenStudy (anonymous):

step by step using L'hopital rule: limit of [1/x+1/cos(π/2+x)] as x approaches 0

OpenStudy (anonymous):

Do you simply want to check your answers or a tutorial on using the rule to answer this question?

OpenStudy (anonymous):

i am stuck in, how???

OpenStudy (anonymous):

can you show me step by step?

OpenStudy (anonymous):

The rule is applied when the said limit becomes infinite at the limit. For example, your equation above can be simplified to \[\lim_{x \rightarrow 0}(\frac{ \sin x - x }{x \sin x })\] --(1) which is \[\frac{ 0 }{ 0 }\] at the limit. Then we apply the rule to limit (1) Do you follow? Rearrange the terms until 0/0 or 0/infinite pops out. then apply the rule.

OpenStudy (anonymous):

huh..

OpenStudy (anonymous):

then you can apply the rule... if 0/0 pops out again, apply again. Good luck on solving it!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome! :)